# Is Less More? Exploring Token Condensation as Training-free Adaptation for CLIP

### v2

@@ Abstract
Contrastive language-image pre-training (CLIP) has shown remarkable generalization ability in image classification. However, CLIP sometimes encounters performance drops on downstream datasets during zero-shot inference. Test-time adaptation methods attempt to mitigate this by adjusting normalization layers, or tuning context prompts with large batch sizes and extensive augmentations, yet these methods are computationally intensive. This raises an important question: Is there a training-free approach that can efficiently address CLIP’s performance drop in such cases? To explore this, we benchmark token condensation techniques, originally designed to enhance the efficiency of vision transformers, on CLIP zero-shot inference tasks. We observe that although token condensation may compromise in-domain accuracy, it surprisingly enhances CLIP’s performance on certain cross-dataset benchmarks. This motivates two key inquiries: (1) Can token condensation serve as a “free-lunch” solution for CLIP zero-shot inference? (2) What criteria should guide condensation—how can essential tokens be identified and redundant ones eliminated? To address these questions, we propose Token Condensation as Adaptation (TCA), a training-free adaptation method for CLIP by pruning class-irrelevant visual tokens while merging class-ambiguous tokens. As the first approach for CLIP’s token efficiency, TCA demonstrates superior performance across cross-dataset tasks, achieving up to a 21.4% improvement over the strongest baseline while reducing GFLOPs by 12.2% to 48.9%, with minimized hyperparameter dependency. Our code is available at here.
@@ 1 Introduction
Online test-time adaptation (TTA) [54] has emerged as a promising strategy to handle distribution shifts encountered during inference [25]. TTA dynamically fine-tunes pretrained models on unlabeled data batches, enhancing generalization by aligning intermediate-layer batch statistics [34], optimizing for first-order flatness in the loss landscape [12], promoting self-supervised consistency across augmentations [66], or tracking model historical weights [24]. Despite the success of traditional TTA methods, they often require computationally expensive tuning of the backbone’s parameters. This challenge is further amplified in CLIP [36], which consist of vast parameter sets and require large batch sizes (e.g., 256) for stabilizing adaptation [8].
@@
Test-time prompting (TPT) offers an efficient alternative for TTA by shifting adaptation focus to the language side of CLIP, learning a small set of task-specific context prompts only for downstream tasks. Nevertheless, TPT largely overlooks the impact of visual distribution shifts. Adapting to high-variance target images through prompts often relies on external source data [42] or extensive data augmentation [11] (e.g., 60$\times$ more AugMix or diffusion-based synthetic samples). In strict online TTA settings, where the batch size is constrained to one, this reliance on augmentation significantly inflates computational costs, leading to a 60$\times$ rise in GFLOPs compared to single-sample processing (i.e., $1108.61$ vs. $17.59$ GFLOPs). The need for gradient backpropagation during inference further increases the computation burden, making exiting TPT suboptimal for many resource-constrained applications.
What if we could address these challenges and adapt the CLIP model efficiently without additional training? In this paper, we take a closer look at whether training-free solutions can achieve this goal. Given that many CLIP visual encoders rely on Vision Transformers (ViTs), we explore ViT acceleration strategies such as token pruning and merging to reduce the number of redundant tokens as not all tokens contribute equally to predictions. While prior studies have often reported that reducing GFLOPs comes at the expense of significant performance degradation in in-distribution validation such as ImageNet-1K, our findings indicate that condensing tokens with lower attentiveness to its <cls> token can not only maintain but even improve CLIP’s performance on certain unseen target datasets (Tab. 1). To deepen our understanding, we conducted a preliminary leave-one-out analysis to investigate the role of individual visual tokens in CLIP’s visual-text alignment (Fig. 2(a)). The empirical results motivate us to target two specific types of tokens for condensation: (1) class-irrelevant background tokens, which may mislead CLIP by emphasizing non-essential regions that differ from the pre-training data; (2) class-ambiguous object tokens, such as cat whiskers or fur, which can overlap across other classes and disperse visual embeddings. However, relying solely on the current sample’s <cls> token for token condensation may not be optimal, as it tends to be overly universal and may not specifically align with the target classes. Although the text encoder in CLIP contains valuable target semantic information, the dimensional discrepancy between textual and visual embeddings restricts direct alignment. To find <cls> tokens that better represent target classes, we trace the historical <cls> tokens from target data streams on per class with the lowest uncertainty across time steps, which we refer to as domain <cls> tokens. As shown in Fig. 2(b), the maintained domain <cls> tokens follow a promising trajectory that gradually aligns with the text embeddings within the same visual space, bridging visual and textual representations, and guiding <cls> attentiveness to be more semantically aligned with target domains.
Building on these insights, we propose Token Condensation as Adaptation (TCA), a domain-aware token condensation strategy to adapt the CLIP model to unseen domains. As shown in Fig. 1, for every time step, the most reliable <cls> tokens are retained as domain <cls> tokens in a domain-aware token reservoir (DTR). These domain <cls> tokens, in turn, act as implicit indicators for token condensation. Specifically, for the current visual sample, we select inattentive tokens by considering both current <cls> tokens and domain <cls> tokens. Then, we utilize consensus across attention heads to retain only the most relevant class-specific patches and merge class-ambiguous tokens by coreset selection algorithm. Empirically, we show that a minimal number of centers is sufficient to stabilize performance, allowing TCA to scale efficiently with linear complexity. Besides, the domain <cls> tokens are explicitly used as token-level classifiers to correct CLIP logits and improve the final visual-text alignment. To our knowledge, this is the first work to explore token condensation in the context of test-time adaptation. The efficiency and effectiveness of the proposed training-free TCA have been rigorously validated across ten fine-grained classification datasets. In extensive evaluations with traditional TTA, prompting, and test-time prompting approaches, our method consistently outperforms the baselines by up to 21.4% compared with the strongest baseline, while reducing GFLOPs cost by 12.2% to 48.9%. For reproducibility, we provide source code in the supplementary material for reference.
@@
Figure 1: An illustration of the proposed Token Condensation as Adaptation (TCA). To adapt CLIP visual embeddings to text embeddings during test-time, TCA utilizes a domain-aware token reservoir (DTR) to retain historical <cls> tokens with the lowest uncertainty as domain <cls>. These domain <cls> assist in (1) condensing tokens with low attentiveness scores (top-right) and (2) acting as token-level classifiers to refine predictions through logit correction, moving visual embeddings $\mathbf{z}_{t}$ toward text embeddings $\mathbf{t}_{c}$.
@@ 2 Related Work
Online Test-time Adaptation. To address performance degradation during test time, online test-time adaptation (TTA) has gained significant attention. Current TTA methods can be categorized into three main types [54]: optimization-based, data-based, and model-based approaches. Optimization-based methods focus on model updates and optimization objectives [14, 31, 66]. A prominent example is Tent [48], which adapts Batch Normalization layers [18] using entropy minimization. SAR [34] extends this approach to Layer Normalization [1] and Group Normalization [59] with sharpness-aware minimization [12]. Data-based methods include augmentations like selective augmentation [51] and adversarial augmentation [47], as well as the use of memory banks [13, 63, 5]. Model-based approaches involve architectural modifications to enhance model adaptability during testing [27, 19, 20, 53]. However, these methods typically depend on large batch sizes and augmentations, which introduce significant latency for online prediction.
@@
Token Condensation in Vision Transformers. Vision transformers have achieved notable success in image recognition tasks, but their deployment is often limited by resource-constrained environments. To address this, various token condensation methods [32, 38, 41, 61, 70, 22, 50] have been proposed to reduce the computational overhead, primarily through two strategies: token pruning and token merging. Token pruning eliminates less informative tokens to save computation, as seen in methods like EViT [26], which retains tokens based on their attentiveness to the <cls> tokens. ATS [9] introduces input-dependent token pruning to adapt to variability across inputs. Token merging, on the other hand, seeks to combine similar tokens to reduce redundancy. For instance, ToME [3] uses bipartite soft matching to merge neighboring tokens that exhibit similarity. Hybrid approaches have also emerged, such as TPS [55], which prunes tokens and transfers information to retained ones using nearest-neighbor matching, and PruMerge [44], which prunes inattentive tokens using interquartile range and merges via k-nearest neighbors. While previous works have focused on enhancing efficiency within pure ViT models, our approach utilizes token condensation from a different perspective: addressing multimodal distribution shifts in VLMs. This shift remains underexplored, particularly in how to use semantic guidance to prune irrelevant visual tokens that introduce ambiguity. By condensing these tokens, we effectively reduce such distribution shifts, enhancing test-time performance while simultaneously lowering computational costs.
3 Our Approach
Problem Set-up. We begin by revisiting online test-time adaptation of CLIP. For a given downstream task $\mathcal{D}_{\operatorname{tar}}$, the test data $\mathbf{x}=\{\mathbf{x}_{t}\}_{t=1}^{T}$ arrives sequentially at each time step $t$. The objective is to adapt CLIP on the fly to classify the incoming test samples into one of $C$ classes, each represented by a textual prompt like “a photo of a <classname>”. CLIP embeds both visual and textual inputs into a shared space. The visual encoder $E_{v}$ extracts visual features $\mathbf{z}_{t}=E_{v}(\mathbf{V}_{t})\in\mathbb{R}^{D}$ from image patches $\mathbf{V}_{t}=[\mathbf{v}_{\operatorname{cls}},\mathbf{v}_{1},\dots,\mathbf{v}_{N}]\in\mathbb{R}^{(N+1)\times D_{v}}$ of dimension $D_{v}$, where $\mathbf{v}_{\operatorname{cls}}$ is a special <cls> token appended to $N$ patches. The text encoder $E_{t}$ generates class embeddings $\mathbf{T}=\{\mathbf{t}_{c}\}_{c=1}^{C}$, where each $\mathbf{t}_{c}\in\mathbb{R}^{D}$ corresponding to a class prompt. Classification is performed by computing the cosine similarity between the visual embedding $\mathbf{z}_{t}$ and each class embedding $\mathbf{t}_{c}$ with the probabilities calculated as:
@@ 3 Our Approach
3.1 Preliminary Studies
Pitsfalls of TTA. Since the target domain $\mathcal{D}_{\operatorname{tar}}$ is unseen during CLIP’s pre-training, the alignment between visual embeddings $\mathbf{z}_{t}$ and the textual embeddings $\mathbf{T}$ may be suboptimal. Previous methods have attempted to address this by learning domain-specific prompts [62] or replacing classifier weights with visual centroids [19] to move $\mathbf{T}$ closer to $\mathbf{z}_{t}$. However, the variability in CLIP’s visual embeddings is often much higher than in textual embeddings [36]. At the patch level, individual tokens within the visual embeddings can drift and vary significantly [36]. Thus, it becomes more urgent to derive methods that adjust $\mathbf{z}_{t}$ towards $\mathbf{T}$ for improved alignment.
(b) Alignment between the updated domain <cls> token and text embedding over time.
Figure 2: Preliminary studies of token influence and the strategy of caching domain <cls> to improve alignment.
How to Mine <cls> Tokens Aligned with Text? In CLIP pretraining, the <cls> token is trained to align with a vast array of concepts, leading to attentiveness patterns that may extend beyond the target classes. Given the mismatch in dimensionality between textual and visual tokens, a fundamental challenge remains in finding a more representative <cls> token that consistently aligns with the text embeddings. To address this, we exploit temporal cues by tracing target samples with the lowest uncertainty as domain <cls>. The similarity between these domain <cls> tokens and $\mathbf{T}$ is plotted in Fig. 2(b). Our findings indicate that retaining these domain <cls> tokens with minimal uncertainty leads to a steady convergence of visual embeddings with the corresponding text embeddings in the same visual space. These domain <cls> tokens serve as contextually relevant proxies, effectively bridging the gap between visual and textual representations and enabling <cls> attentiveness to align more consistently with target classes.
3.2 Token Condensation as Adaptation
Building upon our empirical findings, we propose a novel strategy called Token Condensation as Adaptation (TCA), which selectively removes or merges tokens contributing to drift, enabling efficient adaptation in a training-free manner. Specifically, given an $L$-layer ViT, the forward pass through the $l$-th Transformer block, where $l\in[1,2,\dots,L]$, is formulated as:
@@ 3.2 Token Condensation as Adaptation
where $\mathbf{V}^{l}\in\mathbb{R}^{(N+1)\times D_{v}}$ represents the token embeddings at layer $l$. The matrices $\mathbf{W}^{h}_{Q}$, $\mathbf{W}^{h}_{K}$, $\mathbf{W}^{h}_{V}\in\mathbb{R}^{D_{v}\times D_{v}}$ represent linear projection for the query, key, value vectors in the $h$-th attention head, and $H$ denotes the total number of attention heads. To effectively condense tokens without compromising performance, we introduce two key functions: the token pruning function $f_{\operatorname{prune}}$ and the token merging function $f_{\operatorname{merge}}$. In line with [26], these functions are applied between the multi-head self-attention layers and feed-forward layers of each transformer block. The modified forward pass becomes:
where $f_{\operatorname{prune}}(\cdot;\mathfrak{R}):\mathbb{R}^{N+1}\mapsto\mathbb{R}^{(\alpha\cdot R\cdot N)+1}$ and $f_{\operatorname{merge}}(\cdot;\mathfrak{R}):\mathbb{R}^{(\alpha\cdot R\cdot N)+1}\mapsto\mathbb{R}^{R\cdot N+1}$ are responsible for reducing the number of tokens from $N+1$ (including the <cls> token) to $R\cdot N+1$, where $R$ represents the fraction of tokens to be preserved. The parameter $\alpha$ controls the extent of token pruning, ensuring only the most semantically significant tokens are retained. Detailed explanation of $f_{\operatorname{prune}}$ and $f_{\operatorname{merge}}$ can be found in Sec. 3.3 and Sec. 3.4.

### v3

# Is Less More? Exploring Token Condensation as Training-free Test-time Adaptation

@@ Abstract
Contrastive Language-Image Pretraining (CLIP) excels at learning generalizable image representations but often falls short in zero-shot inference on certain downstream datasets. Test-time adaptation (TTA) mitigates this issue by adjusting components like normalization layers or context prompts, yet it typically requires large batch sizes and extensive augmentations, leading to high computational costs. This raises a key question: Can VLMs’ performance drop in specific test cases be mitigated through efficient, training-free approaches? To explore the solution, we investigate token condensation (TC) techniques, originally designed to enhance vision transformer efficiency by refining token usage during inference. We observe that informative tokens improve visual-text alignment in VLMs like CLIP on unseen datasets. However, existing TC methods often fail to maintain in-distribution performance when reducing tokens, prompting us to ask: How can we transform TC into an effective “free-lunch” adaptation strategy for VLMs? To address this, we propose Token Condensation as Adaptation (TCA), a training-free adaptation method that takes a step beyond standard TC. Rather than passively discarding tokens, TCA condenses token representation by introducing reservoir-based domain anchor tokens for information-preserving token reduction and logits correction. TCA achieves up to a 21.4% performance improvement over the strongest baseline on cross-dataset benchmark and the CIFAR-100-Corrupted dataset while reducing GFLOPs by 12.2% to 48.9%, with minimal hyperparameter dependency on both CLIP and SigLIP series.
@@ 1 Introduction
Online test-time adaptation (TTA) [59] has emerged as a promising strategy to handle distribution shifts encountered during inference [26]. TTA dynamically fine-tunes pretrained models on unlabeled data batches, enhancing generalization by aligning intermediate-layer batch statistics [37], optimizing for first-order flatness in the loss landscape [12], promoting self-supervised consistency across augmentations [72], or tracking model historical weights [25]. Despite the success of traditional TTA methods, they often require computationally expensive tuning of the backbone’s parameters. This challenge is further exacerbated in CLIP [39], SigLIP [70], and SigLIP v2 [51], which rely on visual-text similarity for zero-shot prediction. With vast parameter sets and the need for large batch sizes (e.g., 256) to stabilize adaptation [8], applying conventional TTA to VLMs becomes increasingly impractical.
@@
To circumvent the need for full model tuning, test-time prompting (TPT) has been proposed as a more efficient alternative. By learning a small set of task-specific context prompts, TPT better aligns text features with visual representations, enabling lightweight adaptation. However, TPT primarily focuses on refining text inputs while largely overlooking the impact of visual distribution shifts. Besides, adapting to high-variance target images through prompts often relies on external source data [45] or extensive data augmentation [11] (e.g., 60$\times$ more AugMix or diffusion-based synthetic samples). In strict online TTA settings, where the batch size is constrained to one, this reliance on augmentation significantly inflates computational costs, leading to a 60$\times$ rise in GFLOPs compared to single-sample processing (i.e., $1108.61$ vs. $17.59$ GFLOPs). The need for gradient backpropagation during inference further increases the computation burden, making exiting TPT suboptimal for many resource-constrained applications.
The trade-off between the increased computation of TTA and the efficiency demands of testing raises a key question: Can VLMs achieve TTA while maintaining testing efficiency? In this paper, we explore training-free solutions to address this challenge. Given that many VLM visual encoders rely on Vision Transformers (ViTs), token pruning and merging accelerate inference by reducing redundancy while preserving essential tokens through adjusted token usage in the forward pass. However, efficiency-driven token adjustments often come at the cost of degraded in-distribution performance (e.g., ImageNet-1K validation) in exchange for lower GFLOPs [27]. In contrast, our analysis reveals that selectively condensing low-attentiveness tokens not only preserves performance but can even enhance it on certain unseen datasets (Tab. 1). This insight motivates our exploration of efficient, training-free TTA through token condensation.
To better understand which tokens contribute most to VLM adaptation, we conducted a preliminary analysis on token importance, investigating their role in visual-text alignment (Fig. 2(a)). This analysis highlights two key types of tokens that benefit from condensation: (1) class-irrelevant background tokens, which may mislead the model by emphasizing non-essential regions that deviate from pretraining data distributions, and (2) class-ambiguous object tokens, such as animal fur or textures, which overlap across categories and disperse visual embeddings. However, simply adapting to individual test samples in isolation is suboptimal, as it fails to capture domain-level trends that evolve over time. While textual features encode valuable class-specific information, their embedding space differs from visual representations, limiting direct alignment.
To address these challenges, we introduce Token Condensation as Adaptation (TCA), a training-free adaptation method that dynamically refines token selection based on evolving domain knowledge. Instead of indiscriminately reducing tokens, TCA tracks and maintains domain-representative tokens over time, allowing the model to be adapted to new domains. As shown in Fig. 1, TCA stores domain-aware anchor tokens (e.g., <cls> tokens in CLIP or pooled vectors in SigLIP) in a reservoir (DTR), which then guides adaptation by refining class-specific representations. At each time step, TCA aligns token usage with domain trends by selectively adjusting token attentiveness based on both the current test sample and accumulated domain information in DTR. Additionally, TCA utilizes domain anchor tokens to refine model logits, improving visual-text alignment without modifying model parameters.
To our knowledge, this is the first work to explore token condensation as a form of test-time adaptation. Unlike previous approaches, TCA offers a lightweight, scalable, and training-free solution that generalizes across diverse data domains and can be readily extended to SigLIP and SigLIP v2. The extensive evaluations on the cross-dataset benchmark and the CIFAR-100-Corrupted dataset demonstrate that TCA consistently outperforms traditional TTA, prompting, and test-time prompting methods, achieving up to a 21.4% improvement over the strongest baseline while reducing GFLOPs by 12.2% to 48.9%.
@@
Figure 1: Proposed Token Condensation as Adaptation (TCA). Using CLIP as an example, to adapt visual embeddings to text embeddings during test-time, TCA utilizes a domain-aware token reservoir (DTR) to retain historical <cls> tokens with the lowest uncertainty as domain anchor tokens. These anchor tokens assist in (1) condensing tokens with low attentiveness scores (top-right) and (2) acting as token-level classifiers to refine predictions through logits self-correction, moving visual embeddings $\mathbf{z}_{t}$ toward text embeddings $\mathbf{t}_{c}$.
@@ 2 Related Work
Online Test-time Adaptation. To address performance degradation during test time, online test-time adaptation (TTA) has gained significant attention. Current TTA methods can be categorized into three main types [59]: optimization-, data-, and model-based approaches. Optimization-based methods focus on model updates and optimization objectives [14, 33, 72]. A prominent example is Tent [52], which adapts Batch Normalization layers [19] by entropy minimization. SAR [37] extends this approach to Layer Normalization [1] and Group Normalization [64] with sharpness-aware minimization [12]. Data-based methods include augmentations like selective [55] and adversarial augmentation [50], and memory banks [13, 68, 5]. Model-based approaches involve architectural modifications to enhance model adaptability during testing [29, 20, 21, 58]. However, they typically depend on large batch sizes and augmentations, which introduce significant latency for online prediction.
@@
Token Condensation in Vision Transformers. Vision transformers have achieved notable success in image recognition tasks, but their deployment is often limited by resource-constrained environments. To address this, various token condensation methods [35, 41, 44, 66, 76, 23, 54] have been proposed to reduce the computational overhead, primarily through two strategies: token pruning and token merging. Token pruning eliminates less informative tokens to save computation, as seen in methods like EViT [27], which retains tokens based on their attentiveness to the <cls> tokens. ATS [9] introduces input-dependent token pruning to adapt to variability across inputs. Token merging, on the other hand, seeks to combine similar tokens to reduce redundancy. For instance, ToME [3] uses bipartite soft matching to merge neighboring tokens that exhibit similarity. Hybrid approaches have also emerged, such as TPS [60], which prunes tokens and transfers information to retained ones using nearest-neighbor matching, and PruMerge [47], which prunes inattentive tokens using interquartile range and merges via k-nearest neighbors. While previous works have focused on enhancing efficiency within pure ViT models, our approach goes far beyond a mere adaptation of these methods, addressing multimodal distribution shifts in VLMs. This shift remains underexplored, particularly in how to use semantic guidance to prune irrelevant visual tokens that introduce ambiguity. By condensing these tokens, we effectively reduce such distribution shifts, enhancing test-time performance while simultaneously lowering computational costs (See Tab. 1).
3 Token Condensation as Adaptation
Without loss of generality, we use CLIP as the representative model in the rest of the sections. We show our method can be effortlessly extended to other VLMs in Sec. 4.3.
Problem Set-up. We begin by revisiting online test-time adaptation. For a given downstream task $\mathcal{D}_{\operatorname{tar}}$, the test data $\mathbf{x}=\{\mathbf{x}_{t}\}_{t=1}^{T}$ arrives sequentially at each time step $t$. The objective is to adapt the model on the fly to classify the incoming test samples into one of $C$ classes, each represented by a textual prompt like “a photo of a <classname>”. CLIP embeds both visual and textual inputs into a shared space. The visual encoder $E_{v}$ extracts visual features $\mathbf{z}_{t}=E_{v}(\mathbf{V}_{t})\in\mathbb{R}^{D}$ from image patches $\mathbf{V}_{t}=[\mathbf{v}_{\operatorname{cls}},\mathbf{v}_{1},\dots,\mathbf{v}_{N}]\in\mathbb{R}^{(N+1)\times D_{v}}$ of dimension $D_{v}$, where $\mathbf{v}_{\operatorname{cls}}$ is a CLIP-only <cls> token appended to $N$ patches. The text encoder $E_{t}$ generates class embeddings $\mathbf{T}=\{\mathbf{t}_{c}\}_{c=1}^{C}$, where each $\mathbf{t}_{c}\in\mathbb{R}^{D}$ corresponding to a class prompt. Classification is performed by computing the cosine similarity between the visual embedding $\mathbf{z}_{t}$ and each class embedding $\mathbf{t}_{c}$ with the probabilities calculated as:
@@ 3 Token Condensation as Adaptation
For the visual encoder of ViT-based CLIP, given an $L$-layer ViT, the forward pass through the $l$-th Transformer block, where $l\in[1,2,\dots,L]$, is formulated as:
@@
where $\mathbf{V}^{l}\in\mathbb{R}^{(N+1)\times D_{v}}$ is the token embeddings at layer $l$. The matrices $\mathbf{W}^{h}_{Q}$, $\mathbf{W}^{h}_{K}$, $\mathbf{W}^{h}_{V}\in\mathbb{R}^{D_{v}\times D_{v}}$ are the linear projection matrices for the query, key, and value vectors in the $h$-th attention head, respectively, within the total number of attention heads $H$.
3.1 Empirical Discussion
Pitfalls of TTA. In CLIP, since the target domain $\mathcal{D}_{\operatorname{tar}}$ is unseen during pre-training, the alignment between visual embeddings $\mathbf{z}_{t}$ and the textual embeddings $\mathbf{T}$ may be suboptimal. Previous methods have attempted to address this by learning domain-specific prompts [67] or replacing classifier weights with visual centroids [20] to move $\mathbf{T}$ closer to $\mathbf{z}_{t}$. However, the variability in CLIP’s visual embeddings is often much higher than in textual embeddings [39]. At the patch level, individual tokens within the visual embeddings can drift and vary significantly [39]. Thus, it becomes more urgent to derive methods that adjust $\mathbf{z}_{t}$ towards $\mathbf{T}$ for improved alignment.
(b) Alignment between the updated domain anchor token and text embedding over time.
Figure 2: Empirical studies of token influence and the strategy of caching domain anchor token (i.e., <cls> tokens in CLIP) to improve alignment.
How to Mine <cls> Tokens Aligned with Text? The <cls> token in CLIP visual encoder is trained for broad concept alignment, often extending beyond target classes. Due to the mismatch between textual and visual token spaces, a key challenge is finding a more representative visual <cls> token that better aligns with text embeddings. To address this, we track visual <cls> tokens from test samples with the lowest entropy as domain anchors, updating them dynamically at each time step $t$. We then measure the alignment between these stored <cls> tokens and text embeddings $\mathbf{T}$ over time (Fig. 2(b)). Our results reveal a progressive correlation – as the stored domain anchors consistently favor lower-entropy samples, their <cls> tokens increasingly align with text embeddings, effectively bridging the gap between visual and textual representations.
3.2 Method Overview
Building on our empirical findings, we propose Token Condensation as Adaptation (TCA), a training-free online adaptation strategy that enhances VLMs by filtering out tokens that contribute to visual-text misalignment, ensuring stable and efficient adaptation.
To leverage VLMs’ existing knowledge, TCA introduces a domain-aware token reservoir (DTR) that retains representative anchor tokens for adaptation (Sec. 3.3). Note that in models like SigLIP [51] and SigLIP v2 [51], which lack a <cls> token, we use the pooled feature vector to align with their holistic aggregation strategy. These domain anchors provide a stable reference for adaptation. Guided by these anchors, we perform cross-head token condensation between multi-head self-attention and feed-forward layers [27], selectively merging or discarding less informative tokens to ensure that only relevant, domain-consistent information is retained (Sec. 3.4). Finally, stored domain anchor tokens aid in logits self-correction, refining model predictions based on accumulated domain knowledge (Sec. 3.5).
3.3 Domain-aware Token Reservoir